OpenStudy (anonymous):

The dimensions of a rectangular prism are the same as a Pythagorean triangle with legs of 6 and 8. The volume of the prism is

Directrix (directrix):

The Pythagorean triangle with legs of 6 and 8 has hypotenuse 10. (6,8,10) is a 2-multiple of the triple (3,4,5). Volume = Bh where B is the area of the base. In this case, the base is a rectangle. Volume = L*W*H Volume = 6*8*10 Volume = 480 cubic units.